OpenStudy (anonymous):

MEDAL AND FAN!!!!!!!!! What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)

pooja195 (pooja195):

First find the slope \[\huge~\rm\frac{ y_2-y_1 }{ x_2-x_1}=slope\]

OpenStudy (anonymous):

0

OpenStudy (anonymous):

slope is 3

OpenStudy (anonymous):

what is y -int

OpenStudy (marihelenh):

After you find slope, put it into the equation \[(y-y _{1})=m(x-x _{1})\]

pooja195 (pooja195):

Then we use the point slope formula \[\huge~\rm~ y-y_1=m(x-x_1)\] Where m=slope and then pick whichever point u want to plug in (2, 4) m=3 y1=4 x1=2 from there solve